# Bayesian inference and cure rate modeling for event history data

### v1

@@ B.2 MALA step
$A_{6,h}(\boldsymbol{\theta},\tilde{\boldsymbol{\theta}})=\min\left\{1,\frac{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\tilde{\boldsymbol{\theta}},\boldsymbol{x},\boldsymbol{I})\dot{\pi}_{h}(\tilde{\boldsymbol{\theta}})}{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\boldsymbol{\theta},\boldsymbol{x},\boldsymbol{I})\dot{\pi}_{h}(\boldsymbol{\theta})}\frac{\mathrm{P}\left(\tilde{\boldsymbol{\theta}}\rightarrow\boldsymbol{\theta}\right)}{\mathrm{P}\left(\boldsymbol{\theta}\rightarrow\tilde{\boldsymbol{\theta}}\right)}\right\},$ (24)
@@ D.1 Comparison of Algorithms 1 and 2 on a single simulated dataset
Figure 10: Thinned MCMC outputs for four randomly initialized chains according to Algorithm 1 (left) and Algorithm 2 (right). The horizontal yellow line indicates the values used to generate the dataset.
@@
[…]
@@ D.2 Additional simulation study results
This section contains supplementary Figures 11, 12, 13, 14, 15, 16 and 17, discussed in Section 5.1.
@@
Figure 11: Point estimates (MAP) with $95\%$ Highest Density Intervals for our simulated datasets. Different colour indicate levels of sample size: 500 (), 2000 (), 5000 (). The horizontal line indicates the true value (see Scenarios A1 and A2 in Table 1).
@@
Figure 12: Point estimates (MAP) with $95\%$ Highest Density Intervals for our simulated datasets. Different colour indicate levels of sample size: 500 (), 2000 (), 5000 (). The horizontal line indicates the true value (see Scenarios B1 and B2 in Table 1).
@@
Figure 13: Point estimates (MAP) with $95\%$ Highest Density Intervals for our simulated datasets. Different colour indicate levels of sample size: 500 (), 2000 (), 5000 (). The horizontal line indicates the true value (see Scenarios C1 and C2 in Table 1).
@@
Figure 14: Point estimates (MAP) with $95\%$ Highest Density Intervals for our simulated datasets. Different colour indicate levels of sample size: 500 (), 2000 (), 5000 (). The horizontal line indicates the true value (see Scenarios D1 and D2 in Table 1).
@@
Figure 15: Point estimates (MAP) with $99\%$ Highest Density Intervals for our simulated datasets. Different colour indicate levels of sample size: 500 (), 2000 (), 5000 (). The horizontal line indicates the true value (see Scenarios E1 and E2 in Table 1).
@@
Figure 16: Point estimates (MAP) with $95\%$ Highest Density Intervals for our simulated datasets. Different colour indicate levels of sample size: 500 (), 2000 (), 5000 (). The horizontal line indicates the true value (see Scenarios F1 and F2 in Table 1).
@@
Figure 17: Point estimates (MAP) with $95\%$ Highest Density Intervals for our simulated datasets. Different colour indicate levels of sample size: 500 (), 2000 (), 5000 (). The horizontal line indicates the true value (see Scenarios F3 and F4 in Table 1).

### v2

@@ B.2 MALA step
$A_{6,h}(\boldsymbol{\theta},\tilde{\boldsymbol{\theta}})=\min\left\{1,\frac{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\tilde{\boldsymbol{\theta}},\boldsymbol{x},\boldsymbol{I})\dot{\pi}_{h}(\tilde{\boldsymbol{\theta}})}{\dot{f}_{h}(\boldsymbol{y}|\boldsymbol{\Delta},\boldsymbol{\theta},\boldsymbol{x},\boldsymbol{I})\dot{\pi}_{h}(\boldsymbol{\theta})}\frac{\mathrm{P}\left(\tilde{\boldsymbol{\theta}}\rightarrow\boldsymbol{\theta}\right)}{\mathrm{P}\left(\boldsymbol{\theta}\rightarrow\tilde{\boldsymbol{\theta}}\right)}\right\},$ (B.25)
@@ D.1 Comparison of Algorithms 1 and 2 on a single simulated dataset
The corresponding traces of sampled parameter values are displayed in Figure D.10 (see for example the generated values of $\gamma$ and $\beta_{0}$). In such cases it is quite challenging to properly explore the posterior surface, since the simulated MCMC sample may be trapped in a local mode. In Figure D.10 it is clearly seen that the MC3 implementation (using $C=16$ heated chains) in Algorithm 2 enables the generated parameter values consecutively alternating between the main and the minor ones (this behaviour is vividly illustrated particularly for $\gamma$ and $\beta_{0}$), therefore, the chains are able to freely explore the posterior distribution and avoid getting trapped to the minor mode(s).
We have also checked the validity of our findings in cases of smaller sample sizes. Therefore, we have repeated the present analysis when reducing the size of the simulated data set to $n=300$ observations. We concluded that the MC3 sampler leads to similar answers in four distinct runs under different starting values, as indicated in Figure D.11. On the other hand, when using Algorithm 1 the generated MCMC sample might get trapped into minor modes of the posterior distribution (results not displayed).
@@ D.2 Additional simulation study results
This section contains supplementary Figures D.12, D.13, D.14, D.15, D.16, D.17 and D.18, discussed in Section 5.1.
D.3 Time comparison
In this section we assess the computational complexity of our method by examining the CPU time needed to run our MC3 algorithm for various levels of sample size. We are also reporting the corresponding time needed by the EM algorithm implementation. For this purpose, we generated 20 synthetic datasets as described in the main simulation study, considering that the sample size ($n$) ranges in the set $n\in\{250,500,\ldots,5000\}$. For each dataset we ran the the EM and MCMC algorithms as described in Section 5.1 and measured CPU time. The results are displayed in Figure D.19. Note that for the MCMC implementation we also utilized parallel processing of the heated chains so we are reporting both user and elapsed time. All benchmarks were performed in a Linux workstation with the following specifications: Operating System: Ubuntu 22.04.4 LTS, 11th Gen Intel Core i9-11900 @ 2.50GHz $\times$ 16.
Figure D.10: Thinned MCMC outputs for four randomly initialized chains according to Algorithm 1 (left) and Algorithm 2 (right) for a simulated from Scenario A1 with sample size equal to $n=500$. The horizontal yellow line indicates the values used to generate the dataset.
Figure D.11: Thinned MCMC outputs for four randomly initialized chains according to Algorithm 2 for a simulated from Scenario A1 with sample size equal to $n=300$. The horizontal yellow line indicates the values used to generate the dataset.
@@ D.3 Time comparison
Figure D.12: Point estimates (MAP) with $95\%$ Highest Density Intervals for our simulated datasets. Different colour indicate levels of sample size: 500 (), 2000 (), 5000 (). The horizontal line indicates the true value (see Scenarios A1 and A2 in Table 1).
@@
Figure D.13: Point estimates (MAP) with $95\%$ Highest Density Intervals for our simulated datasets. Different colour indicate levels of sample size: 500 (), 2000 (), 5000 (). The horizontal line indicates the true value (see Scenarios B1 and B2 in Table 1).
@@
Figure D.14: Point estimates (MAP) with $95\%$ Highest Density Intervals for our simulated datasets. Different colour indicate levels of sample size: 500 (), 2000 (), 5000 (). The horizontal line indicates the true value (see Scenarios C1 and C2 in Table 1).
@@
Figure D.15: Point estimates (MAP) with $95\%$ Highest Density Intervals for our simulated datasets. Different colour indicate levels of sample size: 500 (), 2000 (), 5000 (). The horizontal line indicates the true value (see Scenarios D1 and D2 in Table 1).
@@
Figure D.16: Point estimates (MAP) with $99\%$ Highest Density Intervals for our simulated datasets. Different colour indicate levels of sample size: 500 (), 2000 (), 5000 (). The horizontal line indicates the true value (see Scenarios E1 and E2 in Table 1).
@@
Figure D.17: Point estimates (MAP) with $95\%$ Highest Density Intervals for our simulated datasets. Different colour indicate levels of sample size: 500 (), 2000 (), 5000 (). The horizontal line indicates the true value (see Scenarios F1 and F2 in Table 1).
@@
Figure D.18: Point estimates (MAP) with $95\%$ Highest Density Intervals for our simulated datasets. Different colour indicate levels of sample size: 500 (), 2000 (), 5000 (). The horizontal line indicates the true value (see Scenarios F3 and F4 in Table 1).
Figure D.19: Time needed for applying the MCMC and EM algorithms. The elapsed time is smaller than user time for the MCMC implementation because each heated chain runs in parallel cores. Details of algorithms: the EM algorithm uses a random small-EM scheme based on 45 starts and convergence was assessed by iterating the Expectation and Maximization step until the maximum difference between parameter updates in successive steps becomes smaller than 0.005. The MC3 scheme in our MCMC algorithm uses 16 heated chains in total, for a total of $M=20000$ cycles. Each cycle consists of 10 usual MCMC iterations.